OpenStudy (anonymous):

Multiply. (5x)(3x)(6x2 – 1)

OpenStudy (anonymous):

\[15x ^{2}(6x ^{2}-1)\]

OpenStudy (anonymous):

then would it be 90x^4-15x^2 ?

OpenStudy (anonymous):

Yeah but you have to simplify.

OpenStudy (e.cociuba):

(5x)(3x)(6x^2 – 1) 1. work the brackets first 5x * 3x = 15x^2 2. Multiply the above into the left brackets: 15x^2 ( 6x^2 -1) = 15x^2*6x^2 - 15x^2*1 = 90x^4 - 15x^2

OpenStudy (e.cociuba):

Use BODMAS: B Rackets first O OF ( if there is of) D Division ( if there is division) M Multiplication ( if multiplication exists) A Additition ( if there is addition) S Subtraction ( if there is subtraction)

OpenStudy (anonymous):

First, multiply 3x and 5x which is 15x^2 . Then you go ahead and multiply 15x^2 into the (6x^2 -1) which would equal 90x^4 - 15x^2, but usually they make you simplify so you get the GCF (15x^2) and factor it out to simplify.